Convert the parametric equations x = t2 + 1, y = 1 - t to rectangular form.
Answer #1

х 2 t+) Ә) t = 1-9 У= ), + у (1-4)?, 2 1 + + - 2 [a -5) = a*+,-2 а.) 2 1) х с 2 + 1) 1-2 (19) from 2. У 2 (e 2 x - 24 25 а )